OpenStudy (anonymous):

find x and y x (D-1)+Dy=5 Dx+y (D+1)=1

OpenStudy (ankit042):

y= D(1-x) from equation 2 substitute it in first equation to get value of x

OpenStudy (anonymous):

x(D-1) + Dy = 5 Dy = 5 - Dx - x Putting value in second equation Dx + Dy + y = 1 Dx + 5 - Dx - x + (5 - Dx - x)/ D = 1 solve for x then use that value to find y

OpenStudy (anonymous):

Thank you guys, very helpful